Two radioactive nuclides, S and M, both have the same activity X at time t = 0. S and M have a half-life of 26 years and 16 years respectively. Now both samples S and M are mixed. Find out the total activity of the mixture at t = 50 years?


1
Expert's answer
2020-07-13T11:42:59-0400

Let initial number of atom of both S and M is "N_0" .

Since,

"A=\bigg|\frac{dN}{dt}\bigg|=\lambda N"

Thus, initially activity is

"X=\lambda_SN_0+\lambda_MN_0"

Also,

"\lambda=\frac{\ln2}{t_{1/2}}\\
\implies \lambda_S=\frac{\ln2}{26}\\
\lambda_M=\frac{\ln2}{16}"

Hence,

"N_0=\dfrac{208X}{21\ln2}"

Since,

"N=N_0\bigg(\frac{1}{2}\bigg)^{t/t_{\frac{1}{2}}}"

Thus,

"N_S=N_0(1/2)^{50/26}=N_0(1/2)^{25/13}\\N_M=N_0(1/2)^{50/16}=N_0(1/2)^{25/8}"

Thus, activity at t=50 is

"A=\lambda_SN_S+\lambda _MN_M\\
\implies A=\frac{\ln2}{26}N_0(1/2)^{25/13}+\frac{\ln2}{16}N_0(1/2)^{25/8}\\
\implies A=\bigg(\frac{8}{21}(1/2)^{25/13}+\frac{13}{21}(1/2)^{25/8}\bigg)X\approx0.17X"